# Doracamom: Joint 3D Detection and Occupancy Prediction with Multi-view 4D Radars and Cameras for Omnidirectional Perception

### v2

@@ Abstract
3D object detection and occupancy prediction are critical tasks in autonomous driving, attracting significant attention. Despite the potential of recent vision-based methods, they encounter challenges under adverse conditions. Thus, integrating cameras with next-generation 4D imaging radar to achieve unified multi-task perception is highly significant, though research in this domain remains limited. In this paper, we propose Doracamom, the first framework that fuses multi-view cameras and 4D radar for joint 3D object detection and semantic occupancy prediction, enabling comprehensive environmental perception. Specifically, we introduce a novel Coarse Voxel Queries Generator that integrates geometric priors from 4D radar with semantic features from images to initialize voxel queries, establishing a robust foundation for subsequent Transformer-based refinement. To leverage temporal information, we design a Dual-Branch Temporal Encoder that processes multi-modal temporal features in parallel across BEV and voxel spaces, enabling comprehensive spatio-temporal representation learning. Furthermore, we propose a Cross-Modal BEV-Voxel Fusion module that adaptively fuses complementary features through attention mechanisms while employing auxiliary tasks to enhance feature quality. Extensive experiments on the OmniHD-Scenes, View-of-Delft (VoD), and TJ4DRadSet datasets demonstrate that Doracamom achieves state-of-the-art performance in both tasks, establishing new benchmarks for multi-modal 3D perception. Code and models will be publicly available.
Index Terms:
Autonomous driving, camera, 4D radar, deep learning, omnidirectional perception, 3D object detection, 3D occupancy prediction.
@@ I Introduction
Among these sensors, LiDAR operates on the Time of Flight (TOF) principle, emitting and receiving laser beams to generate dense point clouds, providing high-precision geometric representations of the environment [3, 4, 5, 6, 7]. However, LiDAR is susceptible to adverse weather and is costly [8]. In contrast, cameras and radars are more cost-effective, making them suitable for large-scale deployment. Cameras capture rich color and texture information with high resolution but lack depth information and are vulnerable to weather interference [9]. Radar, on the other hand, emits electromagnetic waves to detect target distance, Doppler, and scattering information, offering robustness against weather conditions [10].
The 4D imaging radar, an advancement over conventional radar, not only includes additional elevation information but also provides higher resolution point cloud than conventional $2+1$D radar. Recent studies have shown that it has considerable promise in a various downstream tasks [11, 12]. However, compared to LiDAR, its point cloud remains sparse and noisy. Thus, cross-modal fusion is essential to effectively compensate for these shortcomings, emphasizing the need to integrate information from cameras and 4D radar.
In recent years, especially with the advent of 4D radar datasets, research on 4D radar and camera fusion has demonstrated significant potential in the field of perception. Most mainstream fusion techniques now employ a Bird’s Eye View (BEV) architecture, transforming raw sensor inputs into BEV features for integration. This perspective helps to mitigate occlusion issues with foreground objects and maintains scale consistency. The essence of these methods is the conversion of image data, which lacks depth information, into a bird’s eye view using advanced view transformation algorithms. For occupancy prediction task, most research focuses on vision-centered or vision and LiDAR fusion, as occupancy prediction requires fine-grained voxel representation and semantic information. Traditional radar, lacking height information, is not suitable for 3D occupancy prediction. In contrast, 4D radar offers new possibilities by accessing elevation information and higher resolution point cloud, though related research is still limited. In addition, integrating 3D object detection and occupancy prediction as two key perception tasks within a unified multi-task framework can optimize computational resources and efficiency, providing substantial benefits.
@@ II Related Works
Conventional automotive radar is constrained by its angular resolution, resulting in sparse point clouds lacking height information, necessitating camera-radar fusion for enhanced perception [10]. CenterFusion [38] generates frustum ROIs using 2D detectors and associates them with radar pillars. Feature-level fusion methods work at either BEV level [39, 40] or proposal stage [41, 42]. CRN [39] uses radar for BEV transformation and multi-modal deformable attention for feature alignment. RCBEV [43] combines point fusion and ROI fusion, while RCBEVDet [40] employs RadarBEVNet with dual representations and RCS- aware BEV features, employing a cross-modal attention mech- anism to merge radar and image BEV features. CRAFT [41] associates proposals with radar points in polar coordinates, while CramNet [44] uses ray-constrained attention for geometric correspondence. TransCAR [42] and FUTR3D [45], both extending DETR3D [27], utilize object queries for radar-camera feature interaction.
@@ II-C Recent Progress on Perception with 4D Radar
In the field of 3D object detection, many research efforts have achieved remarkable results by using 4D radar or fusing other sensors. RPFA-Net [64] uses self-attention for 4D radar feature extraction, while SMURF [65] mitigates the sparsity and noise of 4D radar point clouds using pillarization and density features derived from Kernel Density Estimation (KDE) of multi-dimensional Gaussian mixtures. SCKD [66] employs semi-supervised knowledge distillation from LiDAR-radar fusion teacher networks. InterFusion [67] employs a self-attention mechanism to learn features from both LiDAR and 4D radar modalities, exchanging information at intermediate layers. RCFusion [68] leverages RadarPillarNet for hierarchical feature extraction from 4D radar point clouds and effectively fuses BEV features of camera and 4D radar modalities using interactive attention. UniBEVFusion [69] introduces the Radar Depth LSS (RDL) module to improve depth estimation and employs Unified Feature Fusion (UFF) to integrate features from different modalities. LXL [70] proposes a ”radar occupancy-assisted depth-based sampling” strategy that assists image view transformation by combining predicted depth with 3D radar occupancy grids. SGDet3D [71] proposes dual-branch fusion with geometric and semantic information processing, coupled with object-oriented attention for effective radar-camera feature interaction, while HGSFusion [72] introduces hybrid point generation and dual synchronization for radar-camera fusion. Recently, DPFT [73] fuses the 4D radar tensors with camera feature by sampling from queries in 3D space, achieves competitive performance.
For occupancy prediction tasks, RadarOcc [74] achieves binary occupancy prediction using 4D radar tensors, in which only foreground and background are distinguished without providing more specific semantic information. Considering the significant resources consumption required for processing raw radar tensors, such approach is impractical for real-time applications. On the other hand, radar point clouds are widely accepted in ADAS and autonomous driving systems, yet research on 3D semantic occupancy prediction using practical 4D radar point clouds remains limited.
@@ III-A Overall Architecture
The details of the proposed method will be elaborated in the following subsections.
@@ III-B Camera & 4D Radar Encoders
To address the sparsity issue of 4D radar point clouds and obtain their ground velocities by eliminating ego-motion effects, we implement a pre-processing pipeline that combines multi-frame radar point clouds accumulation and velocity compensation. The algorithm processes each radar sweep $s$ using the corresponding ego vehicle velocity $\mathbf{V}_{e}\in\mathbb{R}^{3\times 1}$, which is transformed into the radar coordinate system through the radar-to-ego rotation matrix $\mathbf{R}_{r\to e}\in\mathbb{R}^{3\times 3}$, yielding $\mathbf{V}_{rad}\in\mathbb{R}^{3\times 1}$. To compensate for relative radial velocity, the velocity vector is decomposed into the radial direction based on each point’s azimuth angle $\phi$ and elevation angle $\theta$. The compensated velocity components are then transformed to the current ego coordinate system using rotation matrices $\mathbf{R}_{r\to e}\in\mathbb{R}^{3\times 3}$ and $\mathbf{R}_{e\to e^{\prime}}\in\mathbb{R}^{3\times 3}$. For each point’s position, the transformation is achieved using the radar-to-ego transformation matrices $\mathbf{R}_{r\to e}\in\mathbb{R}^{3\times 3}$ and $\mathbf{t}_{r\to e}\in\mathbb{R}^{3\times 1}$, along with the ego pose transformation matrices from sweep time to current time $\mathbf{R}_{e\to e^{\prime}}\in\mathbb{R}^{3\times 3}$ and $\mathbf{t}_{e\to e^{\prime}}\in\mathbb{R}^{3\times 1}$. Beware the motion of points which introduced by motion of surrounding dynamic objects are omitted during the accumulation operation, since such motions rarely introduce large errors.
The 4D radar encoder processes input point clouds $\mathcal{P}\in\mathbb{R}^{N_{R}\times D}$, where $N_{R}$ and $D$ represent the number of points and attribute dimension, respectively. We adopt RadarPillarNet [68] to encode the input 4D radar point clouds, which generates pseudo images through hierarchical feature extraction. The encoded features are then processed by SECOND and SECONDFPN [77] to produce refined 4D radar BEV features $\mathcal{F}_{R}\in\mathbb{R}^{C_{R}\times H\times W}$, with $C_{R}$ representing the feature dimension and $(H,W)$ denoting the BEV resolution.
@@ III-C Coarse Voxel Queries Generator
We formulate the 3D voxel queries in space as $\mathcal{Q}\in\mathbb{R}^{C\times H_{V}\times W_{V}\times Z_{V}}$, where $(H_{V},W_{V},Z_{V})$ denotes the voxel grid resolution. To reduce computational overhead, we set the BEV plane resolution of the voxel grid as $(H_{V},W_{V})=(\frac{H}{2},\frac{W}{2})$. While existing approaches [32, 35] conventionally utilize random initialization for voxel query generation, this methodology potentially introduces additional complexity to the model training process. To address this limitation and enhance the fidelity of view transformation, we introduce a novel initialization method that integrates geometric priors derived from 4D radar data with semantic features extracted from images. This integration enables the generation of coarse-grained voxel queries with both geometric and semantic priors, establishing a more robust foundation for subsequent refinement procedures. Inspired by [25, 34], we design a voxel query initialization pipeline as followings.
@@
Leveraging $\mathbf{T}_{e\to I}$, we project reference points onto the each image planes to obtain their corresponding coordinates $(x,y,z)$ on the feature maps. Valid points are determined by two criteria: $(x,y)$ must lie within the feature map boundaries and $z$ must be positive. The feature sampling process employs nearest neighbor interpolation, with a ”last-update” strategy resolving overlapping multi-view regions. The final coarse-grained voxel queries are obtained through element-wise addition:
$\mathcal{Q}=\mathcal{Q}_{R}+\mathcal{Q}_{I}$ (3)
@@ III-E Dual-branch Temporal Encoder
Temporal information plays a crucial role in perception systems. Existing methods [21, 78] have demonstrated that leveraging temporal features can effectively address occlusion issues, enhance scene understanding, and improve the accuracy of motion state estimation. However, these approaches are limited to temporal modeling in a single feature space, making it challenging to to capture comprehensive spatio-temporal representations.
@@
where $\mathcal{P}_{t-i}^{3D}$ represents the transformed 3D reference points at historical timestamp $t-i$, and $\mathbf{F}_{(t-i)\rightarrow t}^{vox}\in\mathbb{R}^{C\times H_{V}\times W_{V}\times Z_{V}}$ denotes the aligned historical features, forming a temporal feature sequence $\{\mathbf{F}_{(t-1)\rightarrow t}^{vox},\mathbf{F}_{(t-2)\rightarrow t}^{vox},...,\mathbf{F}_{(t-k)\rightarrow t}^{vox}\}$. Similarly, for the BEV temporal branch, we define 2D reference points $\mathcal{P}_{t}^{2D}\in\mathbb{R}^{H\times W\times 3}$ by setting the height dimension to zero. The historical features are obtained through bilinear interpolation after warping these reference points to corresponding timestamps, which can be formulated as:
@@
where $\mathcal{P}_{t-i}^{2D}$ represents the transformed 2D reference points at historical timestamp $t-i$, and $\mathbf{F}_{(t-i)\rightarrow t}^{BEV}$ denotes the aligned historical features, forming a temporal feature sequence ${\mathbf{F}_{(t-1)\rightarrow t}^{BEV},\mathbf{F}_{(t-2)\rightarrow t}^{BEV},...,\mathbf{F}_{(t-k)\rightarrow t}^{BEV}}$ for the BEV branch.
@@ III-G Multi-task Training
For 3D object detection, following [21], we use DETR-based head unless specified otherwise. The head takes the fused BEV features $\mathbf{F}_{fused}^{BEV}$ as input and directly predicts object categories and attributes. Specifically, each 3D bounding box is parameterized by 10 parameters: dimensions $(l,w,h)$, center location $(x_{o},y_{o},z_{o})$, orientation $(\cos{\theta},\sin{\theta})$, and velocity $(v_{x},v_{y})$. This end-to-end approach eliminates the need for post-processing steps such as NMS. The detection loss $\mathcal{L}_{det}$ consists of a focal loss for classification and an L1 loss for regression, which can be formulated as:
@@
$\mathcal{L}_{aux}=\mathcal{L}_{bce}^{occupied}+\mathcal{L}_{dice}^{occupied}+\mathcal{L}_{bce}^{seg}+\mathcal{L}_{dice}^{seg}$ (19)
@@ IV-A1 Dataset
We evaluate our method on three datasets: OmniHD-Scenes [13], VoD [14], and TJ4DRadSet [15]. All datasets provide data measured by synchronized sensors including 4D imaging radar, camera, and LiDAR, along with 3D object annotations. Notably, OmniHD-Scenes also provides semantic labels for static scenes and dense 3D occupancy ground truth, enabling comprehensive evaluation of our multi-task framework.
The OmniHD-Scenes dataset is a recent large-scale multi-modal dataset featuring a comprehensive omnidirectional sensor suite, which consists of six cameras, six 4D imaging radars, and a 128-beam LiDAR. The dataset contains 1501 sequences with 30-seconds for each, captures under diverse scenarios including nighttime, rainy conditions, and complex traffic situations. Currently, annotations are provided for 200 clips, comprising 3D tracking annotations, static scene segmentation annotations, and dense semantic occupancy ground truth. The 3D bounding box annotations cover four object categories: cars, pedestrians, riders, and large vehicles, while the occupancy ground truth includes 11 semantic classes. The annotated data contains 11921 keyframes in total, with 8321 frames for training and 3600 frames for testing.
The VoD dataset was collected from the campus, suburbs, and old town in Delft, with 5139 training frames and 1296 validation frames. The TJ4DRadSet dataset has 5717 frames for training and 2040 frames for testing, including diverse road scenarios such as urban streets, elevated highways, and industrial areas. Both of them provide synchronized sensor data from a forward-facing setup, consisting of a 4D imaging radar, camera, and LiDAR, along with 3D bounding box annotations. The former contains three categories: cars, pedestrians, and cyclists, while the latter also supplies trucks as supplement.
@@ IV-A2 Evaluation Metrics
For the OmniHD-Scenes dataset, we utilize officially defined metrics to evaluate the performance of 3D detection and occupancy prediction within a detection area of ±60m longitudinally and ±40m laterally around the ego vehicle. For 3D detection, we employ mean Average Precision (mAP) along with four mean True Positive metrics (mTP): mean Average Translation Error (mATE), mean Average Scale Error (mASE), mean Average Orientation Error (mAOE), and mean Average Velocity Error (mAVE). Additionally, we adopt the OmniHD-Scenes Detection Score (ODS) to assess comprehensive performance, as:
$ODS=\frac{1}{8}[4mAP+\sum_{mTP\in TP}(1-\min(1,mTP))]$ (21)
For 3D occupancy prediction evaluation, we employ two key metrics: mean Intersection over Union (mIoU) for semantic accuracy and scene completion IoU for geometric accuracy. The mIoU is calculated by averaging IoU scores across all semantic categories, where IoU measures the overlap between predicted and ground truth occupancy states for each category. The scene completion IoU specifically assesses the model’s capability to distinguish between free and occupied spaces by computing IoU metrics for both spatial states. The details are expressed as follows:
$\mathrm{mIoU}=\frac{1}{C}\sum_{i=1}^{C}\frac{TP_{i}}{TP_{i}+FP_{i}+FN_{i}}$ (22)
$\mathrm{IoU}=\frac{TP}{TP+FP+FN}$ (23)

### v3

@@ Abstract
3D object detection and occupancy prediction are critical tasks in autonomous driving, attracting significant attention. Despite the potential of recent vision-based methods, they encounter challenges under adverse conditions. Thus, integrating cameras with next-generation 4D imaging radar to achieve unified multi-task perception is highly significant, though research in this domain remains limited. In this paper, we propose Doracamom, the first framework that fuses multi-view cameras and 4D radar for joint 3D object detection and semantic occupancy prediction, enabling comprehensive environmental perception. Specifically, we introduce a novel Coarse Voxel Queries Generator that integrates geometric priors from 4D radar with semantic features from images to initialize voxel queries, establishing a robust foundation for subsequent Transformer-based refinement. To leverage temporal information, we design a Dual-Branch Temporal Encoder that processes multi-modal temporal features in parallel across BEV and voxel spaces, enabling comprehensive spatio-temporal representation learning. Furthermore, we propose a Cross-Modal BEV-Voxel Fusion module that adaptively fuses complementary features through attention mechanisms while employing auxiliary tasks to enhance feature quality. Extensive experiments on the OmniHD-Scenes, View-of-Delft (VoD), and TJ4DRadSet datasets demonstrate that Doracamom achieves state-of-the-art performance in both tasks, establishing new benchmarks for multi-modal 3D perception. Code and models can be available at https://github.com/TJRadarLab/Doracamom.
@@ I Introduction
Among these sensors, LiDAR provides high-precision geometric information through dense point clouds [29, 75, 56, 43, 64, 82], but its high cost and weather sensitivity limit deployment [72]. Cameras offer rich visual information but lack depth perception and weather robustness [44]. In contrast, radar provides reliable ranging and velocity measurements in adverse conditions. The all-weather reliability makes it as a key sensor in different fields, like autonomous driving, remote sensing, and SAR images [73, 81, 79]. The 4D imaging radar, advancing beyond conventional 2+12+1D radar, provides additional elevation information and higher resolution point clouds [17, 9]. Therefore, combining 4D radar and cameras brings together the best of both, the all-weather reliability of radar and the rich semantic and texture details from cameras. Compared to expensive LiDAR systems, this fusion scheme offers a more attractive trade-off between performance, robustness, and cost.
Although fusing 4D radar and camera data for perception is a promising research area, existing methods still face several critical challenges. These challenges stem from the heterogeneity between the two modalities, i.e., 4D radar provides sparse coarse-grained geometric priors while cameras provide dense fine-grained semantic features. First, the sparsity of the radar makes query-based methods inefficient as they fail to leverage the valuable geometric priors from the radar. Second, it is a core challenge to effectively fuse these two highly heterogeneous data sources to meet the different feature requirements of multi-task perception. Finally, this combination differs from dense LiDAR-camera fusion, where an “early fusion” before temporal modeling is typically utilized; for sparse 4D radar, this approach is sub-optimal, making it extremely difficult to capture and preserve the unique temporal dynamics of each modality.
@@ II Related Works
Conventional automotive radar is constrained by its angular resolution, resulting in sparse point clouds lacking height information, necessitating camera-radar fusion for enhanced perception [73]. CenterFusion [48] generates frustum ROIs using 2D detectors and associates them with radar pillars. Feature-level fusion methods work at either BEV level [28, 37] or proposal stage [27, 53]. CRN [28] uses radar for BEV transformation and multi-modal deformable attention for feature alignment. RCBEV [87] combines point fusion and ROI fusion, while RCBEVDet [37] employs RadarBEVNet with dual representations and RCS-aware BEV features, employing a cross-modal attention mech anism to merge radar and image BEV features. CRAFT [27] associates proposals with radar points in polar coordinates, while CramNet [25] uses ray-constrained attention for geometric correspondence. TransCAR [53] and FUTR3D [3], both extending DETR3D [61], utilize object queries for radar-camera feature interaction. Notably, RaCFormer [4] further enhances this query-based paradigm by introducing a radar-guided depth estimation to improve the view transformation and adaptively samples features from both PV and BEV views. CRT-Fusion [26] introduces a temporal fusion framework that estimates object motion to guide the alignment and recurrent fusion of multi-timestamp BEV features.
@@ II-C Recent Progress on Perception with 4D Radar
In 3D object detection, many studies have achieved remarkable results using 4D radar and multi-sensor fusion. RPFA-Net [68] and SMURF [40] focus on radar feature extraction, while InterFusion [58] employs a self-attention mechanism to learn features from both LiDAR and 4D radar modalities. RCFusion [84] leverages RadarPillarNet for hierarchical feature extraction from 4D radar point clouds and effectively fuses BEV features using interactive attention. Recent works have introduced more advanced fusion strategies. UniBEVFusion [83] employs RDL module for depth estimation and UFF for multi-modal feature integration. LXL [67] combines predicted depth with radar occupancy grids to enhance image view transformation. SGDet3D [1] utilizes dual-branch fusion and object-oriented attention for feature interaction, while HGSFusion [11] leverages hybrid point generation and dual synchronization. DPFT [10] achieves effective fusion through 3D space query sampling.
For occupancy prediction, RadarOcc [8] performs binary occupancy prediction using 4D radar tensors, but its resource-intensive processing limits real-world applications. While radar point clouds are widely used in autonomous driving systems, research on 3D semantic occupancy prediction with 4D radar point clouds remains limited.
@@ III-B Camera & 4D Radar Encoders
The 4D radar encoder processes input point clouds $\mathcal{P}\in\mathbb{R}^{N_{R}\times D}$, where $N_{R}$ and $D$ represent the number of points and attribute dimension, respectively. Various learning-based methods [80, 78, 68] exist for extracting features from radar data. To balance both efficiency and performance, We adopt RadarPillarNet [84], which generates pseudo images through hierarchical feature extraction. The encoded features are then processed by SECOND and SECONDFPN [70] to produce refined 4D radar BEV features $\mathcal{F}_{R}\in\mathbb{R}^{C_{R}\times H\times W}$, with $C_{R}$ representing the feature dimension and $(H,W)$ denoting the BEV resolution.
@@ III-C Coarse Voxel Queries Generator
We formulate the 3D voxel queries in space as $\mathcal{Q}\in\mathbb{R}^{C\times H_{V}\times W_{V}\times Z_{V}}$, where $(H_{V},W_{V},Z_{V})$ denotes the voxel grid resolution. To reduce computational overhead, we set the BEV plane resolution of the voxel grid as $(H_{V},W_{V})=(\frac{H}{2},\frac{W}{2})$. While existing approaches [63, 62] conventionally utilize random initialization for voxel query generation, this methodology can hinder training efficiency and limit final performance, as the model must learn scene representations without any initial guidance. To address this limitation and enhance the fidelity of view transformation, we introduce a novel initialization method that integrates geometric priors derived from 4D radar data with semantic features extracted from images. This integration enables the generation of coarse-grained voxel queries with both geometric and semantic priors, establishing a more robust foundation for subsequent refinement procedures. Inspired by [18, 20], we design a voxel query initialization pipeline as followings.
@@
Leveraging $\mathbf{T}_{e\to I}$, we project reference points onto the each image planes to obtain their corresponding coordinates $(x,y,z)$ on the feature maps. Valid points are determined by two criteria: $(x,y)$ must lie within the feature map boundaries and $z$ must be positive. The feature sampling process employs nearest neighbor interpolation, with a ”last-update” strategy resolving overlapping multi-view regions. The final coarse-grained voxel queries are obtained by adding $\mathcal{Q}_{R}$ and $\mathcal{Q}_{I}$.
@@ III-E Dual-branch Temporal Encoder
Temporal information plays a crucial role in perception systems. Existing methods [33, 31, 65, 66], have demonstrated that leveraging temporal features can effectively address occlusion issues, enhance scene understanding, and improve the accuracy of motion state estimation. However, these approaches are limited to temporal modeling in a single feature space, making it challenging to to capture comprehensive spatio-temporal representations.
@@
where $\mathcal{P}_{t-i}^{3D}$ represents the transformed 3D reference points at historical timestamp tit-i, and $\mathbf{F}_{(t-i)\rightarrow t}^{vox}\in\mathbb{R}^{C\times H_{V}\times W_{V}\times Z_{V}}$ denotes the aligned historical features, forming a temporal feature sequence $\{\mathbf{F}_{(t-1)\rightarrow t}^{vox},\mathbf{F}_{(t-2)\rightarrow t}^{vox},...,\mathbf{F}_{(t-k)\rightarrow t}^{vox}\}$. Similarly, for the BEV temporal branch, we define 2D reference points $\mathcal{P}_{t}^{2D}\in\mathbb{R}^{H\times W\times 3}$ by setting the height dimension to zero. The historical features are obtained through bilinear interpolation after warping these reference points to corresponding timestamps, which can be formulated as:
@@
where $\mathcal{P}_{t-i}^{2D}$ represents the transformed 2D reference points at historical timestamp tit-i, and $\mathbf{F}_{(t-i)\rightarrow t}^{BEV}$ denotes the aligned historical features, forming a temporal feature sequence $\mathbf{F}_{(t-1)\rightarrow t}^{BEV},\mathbf{F}_{(t-2)\rightarrow t}^{BEV},...,\mathbf{F}_{(t-k)\rightarrow t}^{BEV}$} for the BEV branch.
@@ III-G Multi-task Training
For 3D object detection, following [33], we use DETR-based head unless specified otherwise. The head takes the fused BEV features $\mathbf{F}_{fused}^{BEV}$ as input and directly predicts object categories and attributes. The detection loss $\mathcal{L}_{det}$ consists of a focal loss for classification and an L1 loss for regression, which can be formulated as:
@@
$\mathcal{L}_{aux}=\mathcal{L}_{aux}^{seg}+\mathcal{L}_{aux}^{occupied}$ (18)
@@ IV-A1 Dataset
We evaluate our method on OmniHD-Scenes [86], VoD [51], and TJ4DRadSet [85]. The OmniHD-Scenes features six cameras, six 4D radars, and a 128-beam LiDAR, containing 1501 clips captured in diverse scenarios. The dataset provides annotations for 200 clips, including 3D tracking for four object categories, static scene segmentation, and 11-class semantic occupancy ground truth. The annotated data consists of 11921 keyframes, split into 8321 for training and 3600 for testing.
The VoD dataset covers Delft’s campus, suburbs, and old town, with 5139 training and 1296 validation frames. TJ4DRadSet contains 5717 training and 2040 testing frames from diverse road scenarios. Both datasets provide synchronized data from forward-facing 4D radar, camera, and LiDAR, with 3D bounding box annotations.
@@ IV-A2 Evaluation Metrics
For the OmniHD-Scenes dataset, we utilize officially defined metrics to evaluate the performance of 3D detection and occupancy prediction within a detection area of ±60m longitudinally and ±40m laterally around the ego vehicle. For 3D detection, we employ mean Average Precision (mAP) along with four mean True Positive metrics (mTP): mean Average Translation Error (mATE), mean Average Scale Error (mASE), mean Average Orientation Error (mAOE), and mean Average Velocity Error (mAVE). Additionally, we adopt the OmniHD-Scenes Detection Score (ODS) to assess comprehensive performance. For 3D occupancy prediction evaluation, we employ two key metrics: mean Intersection over Union (mIoU) for semantic accuracy and scene completion IoU for geometric accuracy. The mIoU is calculated by averaging IoU scores across all semantic categories, where IoU measures the overlap between predicted and ground truth occupancy states for each category.